# Mix Q-learning for Lane Changing: A Collaborative Decision-Making Method in Multi-Agent Deep Reinforcement Learning

### v1

@@ IV-B Result of comparative experiments
TABLE VII: The effect of $\varepsilon$ in performance
@@
IV-C Parameter settings of training status
@@
Regarding the latter concern, lane-changing vehicles are homogeneous agents sharing a value network, which causes instability mainly due to the consistency regularisation parameter $\lambda$. In addition to the aforementioned challenges, the reliability of priority settings, which serve as the basis for joint decision making in MQLC, requires further investigation. We explore the rationality of these two types of parameter configurations - the consistency regularising parameter $\lambda$ and the urgency threshold $\varepsilon$ both from experimental results and from the training process. All experiments in this subsection are conducted under the ’normal’ traffic density condition.
IV-C1 The consistency regularisation parameter $\lambda$
As mentioned in chapter III-B1, the setting of the lambda parameter affects the degree to which the individual Q methods influence the global Q, thus affecting both the experimental results and the convergence efficiency of the network. Therefore, its configuration is crucial in the experiments. Table VI and Figure 5(a) show the performance and training process of MQLC models with different $\lambda$ configurations, respectively, while maintaining uniform conditions, under the ’normal’ decision scenario.
@@ IV-C1 The consistency regularisation parameter $\lambda$
TABLE VIII: The results of the ablation experiment
@@
MQLC without $\Omega_{I}$ 29.7 23.48 61.77
@@
IV-C2 The urgency threshold $\varepsilon$
@@
As can be seen from the graph above, more than half of the decision instances show that the urgency $\varepsilon$ for the intelligent agents is less than 0.5. In addition, 9% of the instances have $\varepsilon$ values greater than 1.5, while the remaining instances are almost evenly distributed between the remaining two intervals. We ran tests on the lane-changing performance of models with $\varepsilon$ values of 0.5, 1 and 1.5. The results are shown in table VII and figure 5(b).
@@
(a) Traffic situation(1) at time $t-2$.
(b) Traffic situation(2) at time $t-2$.
(c) Traffic situation(3) at time $t-2$.
(d) Traffic situation(1) at time $t-1$.
(e) Traffic situation(2) at time $t-1$.
(f) Traffic situation(3) at time $t-1$.
@@
Figure 8: The results of the visualisation experiment. The situation (1) and (2) are for MQLC without global decision and the situation (3) is for MQLC. The situation (1) and (2) illustrate two accident scenarios resulting from multi-agent lane-changing decisions without global coordination respectively. In situation (1) at time $t-2$, due to the obstruction caused by vehicles 3 and 4 to intelligent agents 1 and 2, both agents decide to change lanes to the right in order to obtain better driving conditions. At time $t-1$, both agents reach their desired positions, but agent 2 inadvertently obstructs agent 1, forcing it to change lanes to the left, ultimately leading to a collision between vehicles 1 and 3 at time $t$. In situation (2) at time $t-2$, both agents decide to change lanes to the right for better driving conditions. At time $t$, agent 2 rear-ends vehicle 6 as a lack of coordination prevents it from maintaining a safe following distance. Subsequently, agent 1 also rear-ends agent 2, resulting in a serious accident. The situation (3) shows the excellent performance of global decision making in a complex traffic scenario. At time $t-2$, as vehicle 5 changes lanes to the right, driving conditions deteriorate for both agents. Seeking better driving conditions, Agent 2 proactively changes lanes to the left twice, creating space for Agent 1 and improving its own driving visibility. At time $t$, both agents have achieved improved driving conditions.
IV-D Ablation experiment
In order to visually illustrate the specific impact of each innovation on decision making, in this subsection we have conducted ablation experiments on different innovations within MQLC. All experiments in this subsection are conducted under the ’normal’ traffic density condition. The first two components were achieved by removing the relevant parts of the observations and networks from the original model. The third component involved transforming the decision making approach for all agents to be solely determined by individual Q-networks, in order to investigate the impact of the absence of the global decision making component on the experimental results. The final component consisted of configuring all intelligent agents to be controlled solely by the global Q-network, without considering individual agent specifics. These experiments included intention prediction, decision network architecture, global Q involvement, and the concept of decision priorities. The rewards and final results of each model training stage in the experiment are shown in the figure 7 and Table VIII respectively.
@@ IV-D Ablation experiment
he comparison between models without $\Omega_{I}$ and the original model shows that the decision performance of the MQLC model deteriorates significantly when the intention prediction component is missing from the agents’ observations. This degradation is attributed to the fact that intention prediction provides additional useful information to the agents, allowing them to learn the driving patterns of the surrounding vehicles. This allows the agents to implicitly incorporate cooperation with other vehicles into their decisions. This underlines the importance of intention prediction as a fundamental element in achieving collaborative decision making. The comparison between models without $h_{sur}$ and $h_{env}$ and the original model highlights that our improved network, enriched with additional network structures, provides enhanced decision information. This modification proves to be more conducive for the intelligent agents to make rational lane-changing decisions. The MQLC without global decision making, which keeps all other aspects unchanged while adopting the decision making approach of CTDE, initially showed rapid convergence and made the most efficient decisions. However, as training progressed, it gradually fell behind the performance of the original model. This shows that the global decision approach sets a higher bound for the agents, which serves as a basis for decision cooperation. The MQLC model without decision priorities showed the worst performance of all the models compared. The reason for this is that complete reliance on global decision making effectively eliminates consideration of the specific circumstances of individual agents. Consequently, certain agents in disadvantaged positions are forced to adhere to global interests, potentially leading to the formulation of risky decisions. The performance of this model suggests that over-reliance on global coordination-based decision making is not conducive to cooperation among intelligent agents in lane-changing decisions.
IV-E The visualization experiment
@@ V Conclusion and future work
In this paper, to the best knowledge, we propose the first collaborative multi-agent lane-changing decision framework named Mix Q-learning for Lane Changing(MQLC) to improve the performance of lane changing decisions through collaboration of agents. At the individual level, MQLC enhances decision information through intent recognition based on trajectory prediction, adapts the network structure to the characteristics of decision information, and enables more comprehensive feature extraction. At the cooperative level, varying degrees of decision autonomy are granted to the agents according to the urgency of their decisions. Agents make decisions based on individual optimal value functions. Subsequently, the global Q-network coordinates feasible individual decisions using global information to achieve a balance between individual interests and overall benefits. Extensive experimental results on various complex traffic scenarios show that our proposed MQLC model outperforms other state-of-the-art multi-agent lane-changing decision methods in terms of efficiency, safety, and other metrics. The experiments also highlight MQLC’s significant improvement in overall reinforcement learning payoff, and its approach of global coordination of individual decisions has potential applications in other multi-agent reinforcement learning tasks, providing valuable insights into multi-agent collaboration. However, it is worth noting that our collaborative approach focuses primarily on increasing overall reward and does not consider other attributes of vehicle motion, such as speed and comfort. In future work, there is scope to explore more diverse reward criteria to improve the model’s performance on intuitive motion attributes.

### v2

@@ IV-D Design of the intent prediction network
TABLE VIII: The effect of $\varepsilon$ in performance
@@
IV-E Parameter settings of training status
@@
Regarding the latter concern, lane-changing vehicles are homogeneous agents sharing a value network, which causes instability mainly due to the consistency regularisation parameter $\lambda$. In addition to the aforementioned challenges, the reliability of priority settings, which serve as the basis for joint decision making in MQLC, requires further investigation. We explore the rationality of these two types of parameter configurations - the consistency regularising parameter $\lambda$ and the urgency threshold $\varepsilon$ - both from experimental results and from the training process. All experiments in this subsection are conducted under the ’normal’ traffic density condition.
IV-E1 The consistency regularisation parameter $\lambda$
As mentioned in chapter III-B1, the setting of the lambda parameter affects the degree to which the individual Q methods influence the global Q, thus affecting both the experimental results and the convergence efficiency of the network. Therefore, its configuration is crucial in the experiments. Table VII and Figure 6(a) show the performance and training process of MQLC models with different $\lambda$ configurations, respectively, while maintaining uniform conditions, under the ’normal’ decision scenario.
@@ IV-E1 The consistency regularisation parameter $\lambda$
TABLE IX: The results of the ablation experiment
@@
MQLC
without
$\Omega_{I}$ 29.7 23.48 61.77
@@
IV-E2 The urgency threshold $\varepsilon$
@@
As can be seen from the graph above, more than half of the decision instances show that the urgency $\varepsilon$ for the intelligent agents is less than 0.5. In addition, 9% of the instances have $\varepsilon$ values greater than 1.5, while the remaining instances are almost evenly distributed between the remaining two intervals. We ran tests on the lane-changing performance of models with $\varepsilon$ values of 0.5, 1 and 1.5. The results are shown in table VIII and figure 6(b).
@@
(a) Traffic situation(1) at time t2t-2.
(b) Traffic situation(2) at time t2t-2.
(c) Traffic situation(3) at time t2t-2.
(d) Traffic situation(1) at time t1t-1.
(e) Traffic situation(2) at time t1t-1.
(f) Traffic situation(3) at time t1t-1.
@@
Figure 9: The results of the visualisation experiment. The situation (1) and (2) are for MQLC without global decision and the situation (3) is for MQLC. The situation (1) and (2) illustrate two accident scenarios resulting from multi-agent lane-changing decisions without global coordination respectively. In situation (1) at time t2t-2, due to the obstruction caused by vehicles 3 and 4 to intelligent agents 1 and 2, both agents decide to change lanes to the right in order to obtain better driving conditions. At time t1t-1, both agents reach their desired positions, but agent 2 inadvertently obstructs agent 1, forcing it to change lanes to the left, ultimately leading to a collision between vehicles 1 and 3 at time $t$. In situation (2) at time t2t-2, both agents decide to change lanes to the right for better driving conditions. At time $t$, agent 2 rear-ends vehicle 6 as a lack of coordination prevents it from maintaining a safe following distance. Subsequently, agent 1 also rear-ends agent 2, resulting in a serious accident. The situation (3) shows the excellent performance of global decision making in a complex traffic scenario. At time t2t-2, as vehicle 5 changes lanes to the right, driving conditions deteriorate for both agents. Seeking better driving conditions, Agent 2 proactively changes lanes to the left twice, creating space for Agent 1 and improving its own driving visibility. At time $t$, both agents have achieved improved driving conditions.
IV-F Ablation experiment
In order to visually illustrate the specific impact of each innovation on decision making, in this subsection we have conducted ablation experiments on different innovations within MQLC. All experiments in this subsection are conducted under the ’normal’ traffic density condition. The first two components were achieved by removing the relevant parts of the observations and networks from the original model. The third component involved transforming the decision making approach for all agents to be solely determined by individual Q-networks, in order to investigate the impact of the absence of the global decision making component on the experimental results. The final component consisted of configuring all intelligent agents to be controlled solely by the global Q-network, without considering individual agent specifics. These experiments included intention prediction, decision network architecture, global Q involvement, and the concept of decision priorities. The rewards and final results of each model training stage in the experiment are shown in the figure 8 and Table IX respectively.
@@ IV-F Ablation experiment
The comparison between models without $\Omega_{I}$ and the original model shows that the decision performance of the MQLC model deteriorates significantly when the intention prediction component is missing from the agents’ observations. This degradation is attributed to the fact that intention prediction provides additional useful information to the agents, allowing them to learn the driving patterns of the surrounding vehicles. This allows the agents to implicitly incorporate cooperation with other vehicles into their decisions. This underlines the importance of intention prediction as a fundamental element in achieving collaborative decision making. The comparison between models without $h_{sur}$ and $h_{env}$ and the original model highlights that our improved network, enriched with additional network structures, provides enhanced decision information. This modification proves to be more conducive for the intelligent agents to make rational lane-changing decisions. The MQLC without global decision making, which keeps all other aspects unchanged while adopting the decision making approach of CTDE, initially showed rapid convergence and made the most efficient decisions. However, as training progressed, it gradually fell behind the performance of the original model. This shows that the global decision approach sets a higher bound for the agents, which serves as a basis for decision cooperation. The MQLC model without decision priorities showed the worst performance of all the models compared. The reason for this is that complete reliance on global decision making effectively eliminates consideration of the specific circumstances of individual agents. Consequently, certain agents in disadvantaged positions are forced to adhere to global interests, potentially leading to the formulation of risky decisions. The performance of this model suggests that over-reliance on global coordination-based decision making is not conducive to cooperation among intelligent agents in lane-changing decisions.
IV-G The visualization experiment
@@ V Conclusion and future work
In this paper, to the best knowledge, we propose the first collaborative multi-agent lane-changing decision framework named Mix Q-learning for Lane Changing (MQLC) to improve the performance of lane changing decisions through collaboration of agents. At the individual level, MQLC enhances decision information through intent recognition based on trajectory prediction, adapts the network structure to the characteristics of decision information, and enables more comprehensive feature extraction. At the cooperative level, varying degrees of decision autonomy are granted to the agents according to the urgency of their decisions. Agents make decisions based on individual optimal value functions. Subsequently, the global Q-network coordinates feasible individual decisions using global information to achieve a balance between individual interests and overall benefits. Extensive experimental results on various complex traffic scenarios show that our proposed MQLC model outperforms other state-of-the-art multi-agent lane-changing decision methods in terms of efficiency, safety, and other metrics. The experiments also highlight MQLC’s approach of global coordination of individual decisions has potential applications in other multi-agent reinforcement learning tasks, providing valuable insights into multi-agent collaboration. However, it is worth noting that our collaborative approach focuses primarily on increasing overall reward and does not consider other attributes of vehicle motion, such as speed and comfort. In future work, there is scope to explore more diverse reward criteria to improve the model’s performance on intuitive motion attributes.